Find the second derivative of y = (x2+x+1)2


1
Expert's answer
2022-03-03T11:43:38-0500

Solution;

Given;

"y=(x^2+x+1)^2"

"\frac{dy}{dx}=2(x^2+x+1)\frac{d}{dx}(x^2+x+1)"

"y'=2(x^2+x+1)(2x+1)"

Apply product rule;

"y''=2(x^2+x+1)\frac{d}{dx}(2x+1)+2(2x+1)\frac{d}{dx}(x^2+x+1)"

"y''=4(x^2+x+1)+2(2x+1)(2x+1)"

"y''=4(x^2+x+1)+2(2x+1)^2"